OpenStudy (anonymous):

If both the numerator and denominator of 48/64 are divided by 8, the fraction will be in lowest terms. True False i cant do fractions. :(

OpenStudy (donblue22):

True

OpenStudy (anonymous):

the fraction is \[\frac{ 48 }{ 64 }\div \frac{ 8 }{ 8 }\]

OpenStudy (anonymous):

\[=\frac{ 6 }{ 8 }\]

OpenStudy (anonymous):

so false

OpenStudy (donblue22):

Sorry miscalculated

OpenStudy (anonymous):

uhmmm okay lol i get it.

OpenStudy (anonymous):

the lowest terms would be 3/4

OpenStudy (donblue22):

somehow got 5 instead og 6

OpenStudy (anonymous):

ok